OpenStudy (abbycross167):

I WILL FAN AND GIVE A MEDAL!! Can someone please help me? My teacher said that need to redo this problem and I did but I'm stuck on a step... Please help me, I really need it

OpenStudy (abbycross167):

Question~ Find the point-slope form of the equation that goes through these two points. (1, 20) and (8, 4.5)

OpenStudy (abbycross167):

My Answer~ Y= y2-y1/x2-x1 (4.5-20)/(8-1) = -15.5/7 Rewrite -15.5/7 as (-15.5)/7 Rewrite -15.5 as -31/2 = (-31/2) (7/1) Divide: (-31/2) / (7/1) = -31/14

OpenStudy (abbycross167):

I don't know what to do after -31/14

OpenStudy (trojanpoem):

y = mx + b m = slope ( plug the slope in the equation) use any of the points given and solve for b

OpenStudy (abbycross167):

Could you sow me how to do that? Please?

OpenStudy (abbycross167):

@TrojanPoem ?

OpenStudy (trojanpoem):

y = -31/14 x + b (1,20) << (x,y) 20 = -31/14 + b b = 20 + 31/14 20 * 14/14 = 280/14 311/14 = b y = -31/14 x + 311/14

OpenStudy (abbycross167):

So after the -31/14 I add what you just said?

OpenStudy (abbycross167):

@TrojanPoem

OpenStudy (trojanpoem):

As the two points lie in the line, you can let x, y be them and then find b

OpenStudy (abbycross167):

@TrojanPoem Sir/ Ma'am I don't understand what you just said

OpenStudy (trojanpoem):

y = mx + b , you have to find an equation like this to solve the equation.